### v1

# Congruences for the ratios of Rankin–Selberg $L$-functions

Abstract.
A well-known principle states that a congruence between objects should give rise to a corresponding congruence between the special values of $L$-functions attached to these objects. In this article we computationally investigate this principle for Rankin–Selberg $L$-functions attached to pairs of holomorphic cuspforms.
2020 Mathematics Subject Classification:
11F33; 11F30, 11F67
1. Introduction
A well-known principle with origins in Iwasawa theory states that a congruence between objects should translate to a congruence between the special values of $L$-functions attached to these objects. The reader is referred to Vatsal [19] for an instance of this principle, and for a brief discussion of its historical origins. In this article we computationally investigate this principle for Rankin–Selberg $L$-functions attached to pairs of holomorphic cuspforms.
Suppose we have two primitive holomorphic cusp forms $f,f^{\prime}\in S_{k}(\Gamma_{1}(N))$; by a form being primitive, one means it is an eigenform, a newform, and is normalised by taking the first Fourier coefficient to be $1$. Suppose $f(z)=\sum_{n=1}^{\infty}a(n,f)e^{2\pi inz}$ and $f^{\prime}(z)=\sum_{n=1}^{\infty}a(n,f^{\prime})e^{2\pi inz}$ are their Fourier expansions. Let $\mathbb{Q}(f)$ (resp., $\mathbb{Q}(f^{\prime})$) denote the number field generated by the Fourier coefficients of $f$ (resp., $f^{\prime}$), and take $F$ to be any number field containing $\mathbb{Q}(f)$ and $\mathbb{Q}(f^{\prime})$. Let $\mathfrak{P}$ be a prime ideal of the ring of integers of $F$. We say $f$ is congruent to $f^{\prime}$ modulo $\mathfrak{P}$ if $a(n,f)\equiv a(n,f^{\prime})\pmod{\mathfrak{P}}$ for all $n\geq 1$; denote this congruence as $f\equiv f^{\prime}\pmod{\mathfrak{P}}$.
Suppose $g\in S_{l}(\Gamma_{1}(M))$. For the Rankin–Selberg $L$-function $L(s,f\times g)$ to have critical points one needs $k\neq l$. Assume for the moment that $k>l$. Then the set of critical points of $L(s,f\times g)$ is $\{m\in\mathbb{Z}:l\leq m\leq k-1\}$. Since $f$ and $f^{\prime}$ have the same weight, this set is also the set of critical points for $L(s,f^{\prime}\times g).$ For any of these critical points, one expects a congruence between the algebraic parts of $L(m,f\times g)$ and $L(m,f^{\prime}\times g).$ Of course, one has to be careful in clarifying what one means by the algebraic part of an $L$-value. Keeping in mind the shape of the rationality results due to Shimura for these $L$-values (see Thm. 2.7), one observes that if one had at least two critical points, then the ratio $L(m,f\times g)/L(m+1,f\times g)$ of successive critical values for the completed $L$-function is algebraic and lies in the compositum of the number fields $\mathbb{Q}(f)$ and $\mathbb{Q}(g)$. Let us enlarge the number field $F$ above also to contain $\mathbb{Q}(g)$. The purpose of this paper is to investigate the principle:
$f\equiv f^{\prime}\pmod{\mathfrak{P}}\ \ \implies\ \ \frac{L(m,f\times g)}{L(m+1,f\times g)}\equiv\frac{L(m,f^{\prime}\times g)}{L(m+1,f^{\prime}\times g)}\pmod{\mathfrak{P}}$.
It is possible that the ratios of $L$-vaues are not integral in $F$; the congruence is to interpreted as the difference of the ratios has a positive $\mathfrak{P}$-adic valuation.
An obvious variation to consider is to take $k<l,$ i.e., freeze the larger weight cuspform $f$ and take two congruent cuspforms $g,g^{\prime}$ of smaller weight. One may also consider a congruence between a cuspform and an Eisenstein series such as the celebrated Ramanujan congruence. If $g$ is an Eisenstein series then $L(s,f\times g)$ is a product of the form $L(s,f\otimes\chi_{1})\cdot L(s,f\otimes\chi_{1})$ for algebraic Hecke characters $\chi_{1}$ and $\chi_{2}$ of $\mathbb{Q}$.
In this article, we verify that the expected congruences hold in the following cases:
  1. (1)
    $f,f^{\prime}\in S_{24}(\mathrm{SL}_{2}(\mathbb{Z}))$, $\mathfrak{P}=144169$, $g\in S_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$.
  2. (2)
    $f,f^{\prime}\in S_{30}(\mathrm{SL}_{2}(\mathbb{Z}))$, $\mathfrak{P}=51349$, $g\in S_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$.
  3. (3)
    $f,f^{\prime}\in S_{13}(\Gamma_{0}(3),\chi)$, $\mathfrak{P}=13$, $g\in S_{6}(\Gamma_{0}(3))$.
  4. (4)
    $f\in S_{26}(\mathrm{SL}_{2}(\mathbb{Z}))$, and $g,g^{\prime}\in S_{13}(\Gamma_{0}(3),\chi)$, $\mathfrak{P}=13$.
  5. (5)
    $f\in S_{24}(\mathrm{SL}_{2}(\mathbb{Z}))$, $\in S_{12}(\mathrm{SL}_{2}(\mathbb{Z})), gg^{\prime}$ a weight $12$ Eisenstein series, $\mathfrak{P}=691.$
    (This is the Ramanujan congruence.)
In (1) and (2), $f^{\prime}$ is a Galois conjugate of $f$, but in (3) $f^{\prime}$ is not a Galois conjugate of $f$.
The computations rely on two algorithms. The first algorithm, reviewed in 2.2, computes the special values of Rankin-Selberg $L$-functions, and is essentially due to Shimura and Hida. Roughly speaking, it relies on the interpretation of the $L$-value as a Petersson inner product of a cusp form with a holomorphic projection which is due to Shimura [15], and the computation of the holomorphic projection recursively from the given data and an appropriate Eisenstein series due to Hida [5]. The second well-known algorithm, reviewed in 2.3, computes the $L$-values of modular forms from the Hecke-equivariant pairing between a space of cusp forms and a suitable space of modular symbols. It should be mentioned that there is also an algorithm by Tim Dokchitser [2] to compute the $L$-values of modular forms numerically. All the computations in this paper are done on SAGE [14].
Based on these examples, we formulate a conjecture in Conj. 4.1. In the companion article [11], we prove a variation of this conjecture using the machinery of Eisenstein cohomology as developed by Harder and the second author in [3]. As the proofs in [3] involves a cohomological interpretation of Langlands’s theorem on the constant term of an Eisenstein series, the results and proofs in [11] are completely independent of the computational techniques this paper.
2. Preliminaries and the main algorithms
In this section, we begin by reviewing some preliminaries from Shimura [15] on modular forms and the special values of Rankin-Selberg $L$-functions and some results of Hida [5] on the holomorphic projection. We present two algorithms which will be used to verify the congruences mentioned in the introduction.
2.1. Review of some results of Shimura and Hida
Let $\Gamma_{1}(N)$ be the standard Hecke congruence subgroup of $\mathrm{SL}_{2}(\mathbb{Z})$; the volume of its fundamental domain $\Phi_{1}(N)$ with respect to the measure $y^{-2}dxdy$ will be denoted $\mu(\Phi_{1}(N)).$ Let $M_{k}(\Gamma_{1}(N))$ be the space of holomorphic modular forms of weight $k$ for $\Gamma_{1}(N)$. For $f,\,h\in M_{k}(\Gamma_{1}(N))$ such that $fh$ is a cusp form, define the Petersson inner product $\langle f,h\rangle$ by:
$\langle f,h\rangle\ :=\ \mu(\Phi_{1}(N))^{-1}\int_{\Phi_{1}(N)}\overline{f(z)}h(z)y^{k-2}\,dxdy.$
Let $\chi$ be a Dirichlet character modulo $N$. Let $M_{k}(N,\chi)$ (resp., $S_{k}(N,\chi)$) denote the space of holomorphic modular (resp., cusp) forms of weight $k$ for $\Gamma_{0}(N)$ with nebentypus character $\chi$. Assume $l<k$. Suppose $f=\sum_{n=1}^{\infty}a(n,f)q^{n}\in S_{k}(N,\chi)$ and $g=\sum_{n=0}^{\infty}a(n,g)q^{n}\in M_{l}(N,\psi),$ where $q=e^{2\pi iz}$, then the Rankin–Selberg convolution is defined as the Dirichlet series:
$D(s,f\times g)\ :=\ \sum_{n=1}^{\infty}a(n,f)a(n,g)n^{-s}.$
For an integer $\lambda\geq 0$ and a Dirichlet character $\omega$ modulo $N$ such that $\omega(-1)=(-1)^{\lambda}$, define an Eisenstein series:
$E^{*}_{\lambda,N}(z,s,\omega)$ $=\sum_{\gamma\in\Gamma_{\infty}\backslash\Gamma_{0}(N)}\omega(d)(cz+d)^{-\lambda}|cz+d|^{-2s},\quad(\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix})$,
where $\Gamma_{\infty}=\bigl\{\pm\begin{pmatrix}1&m\\ 0&1\end{pmatrix}|m\in\mathbb{Z}\bigr\}$. For the same $\lambda$ and $\omega,$ define:
$E_{\lambda,N}(z,s,\omega)=\sum_{(0,0)\neq(m,n)\in\mathbb{Z}^{2}}\omega(n)(mNz+n)^{-\lambda}|mNz+n|^{-2s}.$
The following relation holds between $E_{\lambda,N}^{*}$ and $E_{\lambda,N}$:
$E_{\lambda,N}(z,s,\omega)=2L_{N}(2s+\lambda,\omega)E^{*}_{\lambda,N}(z,s,\omega),$
where $L_{N}(s,\omega)$ denotes the Dirichlet $L$-series $\sum_{n=1}^{\infty}\omega(n)/n^{s}$ and the subscript $N$ means that with $\omega(n)=0$ if ${\rm gcd}(n,N)\neq 1.$ If $\omega$ is primitive modulo $N$ then this condition is automatic and the subscript may be dropped. One has the following integral representation:
$2(4\pi)^{-s}\Gamma(s)L_{N}(2s+2-k-l,\chi\psi)D(s,f\times g)\ =\\ \int_{\Phi_{0}(N)}\bar{f_{\rho}}g\cdot E_{k-l,N}(z,s+1-k,\chi\psi)\,y^{s-1}\,dxdy.$
Here $\Phi_{0}(N)$ is the fundamental domain for $\Gamma_{0}(N)\backslash\mathbb{H}$ and $f_{\rho}(z)=\overline{f(-\overline{z})}=\sum_{n=1}^{\infty}\overline{a_{n}}e(nz)$. Define $E_{\lambda,N}^{*}(z,\omega)=E_{\lambda,N}^{*}(z,0,\omega)$.
The Maaß–Shimura differential operators on the space of smooth functions on the Poincare upper half plane $\mathbb{H}$ by
$\delta_{\lambda}=\frac{1}{2\pi i}\left(\frac{\lambda}{2iy}+\frac{\partial}{\partial z}\right),\ \ 0<\lambda\in\mathbb{Z},\quad{\rm and}\quad\delta_{\lambda}^{(r)}=\delta_{\lambda+2r-2}\cdots\delta_{\lambda+2}\delta_{\lambda},\ \ 0\leq r\in\mathbb{Z}.$
It is understood that $\delta_{\lambda}^{(0)}$ is the identity operator. Define $d=\frac{1}{2\pi i}\frac{\partial}{\partial z}=q\frac{d}{dq}.$ Then
(1) $\delta_{\lambda}^{(r)}\ =\ \sum_{0\leq t\leq r}\binom{r}{t}\frac{\Gamma(\lambda+r)}{\Gamma(\lambda+t)}(-4\pi y)^{t-r}d^{t},$
where the symbol $d^{0}$ means the identity operator; see [15]. The action of these operators on $1/4\pi y$ is given by the following
Lemma 2.1.
$d^{n}\left(\frac{1}{4\pi y}\right)=\frac{n!}{(4\pi y)^{n+1}},\quad n=0,1,2,\dots,$
$\delta_{\lambda}^{(r)}\left(\frac{1}{4\pi y}\right)=(-4\pi y)^{-(r+1)}\cdot\sum_{t=0}^{r}\binom{r}{t}(-1)^{t+1}\frac{\Gamma(\lambda+r)\Gamma(t+1)}{\Gamma(\lambda+t)}.$
Proof.
The first is proved by induction on $n$; the second from the first and (1). ∎
Let $A_{r}$ denote the set of all functions of the form $h(z)=\sum_{\nu=0}^{r}y^{-nu}g_{\nu}(z)$ with holomorphic functions $g_{\nu}$ on $\mathbb{H}$ which have Fourier expansions $g_{\nu}(z)=\sum_{n=0}^{\infty}b_{\nu n}q^{n}$. Elements of $A_{r}$ may be called nearly holomorphic modular forms. The following lemma is [15, Lem. 7].
Lemma 2.2.
For a positive integer $k>2r$ and a Dirichlet $\chi$ modulo $N$, suppose that an element $h$ of $A_{r}$ satisfies
(2) $\begin{split}h(\gamma(z))(cz+d)^{-k}\in A_{r}\>\>\mbox{for all}\>\>\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in SL_{2}(\mathbb{Z})\\ h(\gamma(z))(cz+d)^{-k}=\chi(d)h(z)\>\>\mbox{for all}\>\>\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma_{0}(N).\end{split}$
Then $h(z)=\sum_{\nu=0}^{r}\delta_{k-2\nu}^{(\nu)}h_{\nu},$ with $h_{\nu}$ of $M_{k-2\nu}(N,\chi)$ being uniquley determined by $h$.
The following orthogonality relation is [15, Lem. 6].
Lemma 2.3.
Suppose $f\in S_{k}(N,\chi)$, $g\in M_{l}(N,\bar{\chi}),$ and $k=l+2r$ with a positive integer $r$. Then $\langle f_{\rho},\delta_{l}^{(r)}g\rangle=0$.
As a consequence of the above two lemmas one has:
Lemma 2.4.
$\langle f_{\rho},\delta_{\lambda}^{(r)}E_{\lambda,N}\rangle\ =\ \langle f_{\rho},h_{0}\rangle.$
The following theorem is one of the main results from [15]:
Theorem 2.5.
Suppose $f\in S_{k}(N,\chi),$ $g\in M_{l}(N,\psi)$, and $l+2r<k$ with a non-negative integer $r$. Then
$D(k-1-r,f\times g)\ =\ c_{r}\pi^{k}\langle f_{\rho},\ g.\delta_{\lambda}^{(r)}E_{\lambda,N}^{*}(z,\chi\psi)\rangle,$
where $\lambda=k-l-2r$, and $c_{r}$ is the rational number explicitly given by:
$c_{r}\ =\ \frac{\Gamma(k-l-2r)}{\Gamma(k-1-r)\Gamma(k-l-r)}.\frac{(-1)^{r}4^{k-1}N}{3}\prod_{p|N}(1+p^{-1}),$
where the product is taken over all prime factors $p$ of $N$.
When $\lambda=2$ and $\omega=1\!\!1$ is the trivial character modulo $N$, then $E_{\lambda,N}^{*}(z,\omega)$ is not holomorphic at $\infty$; its Fourier expansion at $\infty$ is given by
$E_{2,N}^{*}(z,1\!\!1)\ =\ \frac{c}{4\pi y}+\sum_{n=0}^{\infty}c_{n}q^{n},$
where $c$ and all the $c_{n}$ are in $\mathbb{Q}$. This is due to Hecke; see, for example, Miyake [9, p. 288].
Let $f\in S_{k}(N,\chi)$ and $g\in M_{k}(N,\psi)$. Then as noted above
$E_{\lambda,N}^{*}(z,\chi\psi)\ =\ \frac{c}{4\pi y}+\sum_{n=0}^{\infty}c_{n}q^{n}\ =:\ \frac{c}{4\pi y}+E,$
with $c\neq 0$ only when $\lambda=2$ and $\chi\psi=1\!\!1$. Then, by Lemma 2.1 and Lemma 2.2
(3) $g.\delta_{\lambda}^{(r)}E_{\lambda,N}^{*}=g\cdot\delta_{\lambda}^{(r)}\left(\frac{c}{4\pi y}+E\right)\\ =g\cdot\frac{c\cdot c^{\prime}}{(-4\pi y)^{r+1}}+g\cdot\delta_{\lambda}^{(r)}(E)=\begin{cases}\sum_{\nu=0}^{r+1}\delta_{k-2\nu}^{(\nu)}h_{\nu}&\mbox{if $\chi\psi=1\!\!1,\lambda=2$,}\\ \sum_{\nu=0}^{r}\delta_{k-2\nu}^{(\nu)}h_{\nu}&\text{otherwise},\end{cases}$
where $h_{\nu}\in M_{k-2\nu}(N,\chi\cdot\psi)$ and $c^{\prime}\in\mathbb{Q}$ is obtained from Lemma 2.1.
Lemma 2.6.
The $h_{\nu}$, as in equation (3), satisfy the following recursive relations:
  1. (i)
    When $\lambda=2$ and $\chi\psi=\mbox{trivial}$ then
    $\frac{\Gamma(k-(r+1))}{\Gamma(k-2(r+1))}h_{r+1}=cg\sum_{t=0}^{r}\binom{r}{t}(-1)^{t+1}\frac{\Gamma(\lambda+r)\Gamma(t+1)}{\Gamma(\lambda+t)}$
    and
    $\frac{\Gamma(k-\nu)}{\Gamma(k-2\nu)}h_{\nu}=\binom{r}{r-\nu}\frac{\Gamma(\lambda+r)}{\Gamma(\lambda+r-\nu)}gd^{r-\nu}E_{\lambda,N}^{*}(z,\chi\psi)\\ -\sum_{n=\nu+1}^{r+1}\binom{n}{n-\nu}\frac{\Gamma(k-n)}{\Gamma(k-n-\nu)}d^{n-j}h_{n}\qquad\text{for $\nu=0,1,\dots,r$}.$
  2. (ii)
    When $\lambda\neq 2$ or $\chi\psi\neq\mbox{trivial}$ then
    $\frac{\Gamma(k-\nu)}{\Gamma(k-2\nu)}h_{\nu}=\binom{r}{r-\nu}\frac{\Gamma(\lambda+r)}{\Gamma(\lambda+r-\nu)}gd^{r-\nu}E_{\lambda,N}^{*}(z,\chi\psi)\\ -\sum_{n=\nu+1}^{r}\binom{n}{n-\nu}\frac{\Gamma(k-n)}{\Gamma(k-n-\nu)}d^{n-j}h_{n}\qquad\text{for $\nu=0,1,\dots,r$}.$
Proof.
Statement $(ii)$ is proved in Hida [5, p.177]. For $(i)$, when $\lambda=2$ and $\chi\psi=\mbox{trivial}$,
$g.\delta_{\lambda}^{(r)}E_{\lambda,N}^{*}=g\cdot\delta_{\lambda}^{(r)}\left(\frac{c}{4\pi y}+E\right)=g\cdot\delta_{\lambda}^{(r)}\left(\frac{c}{4\pi y}\right)+g\cdot\delta_{\lambda}^{(r)}\left(E\right)=\sum_{n=0}^{r+1}\delta_{k-2n}^{(n)}h_{n}.$
Use (1) and expand both sides of the above equation to get:
$g\cdot\sum_{t=0}^{r}\binom{r}{t}\frac{\Gamma(\lambda+r)}{\Gamma(\lambda+t)}(-4\pi y)^{t-r}d^{t}\left(\frac{c}{4\pi y}+E\right)\\ =\sum_{n=0}^{r+1}\sum_{t=0}^{n}\binom{n}{t}\frac{\Gamma(k-n)}{\Gamma(k-2n+t)}(-4\pi y)^{t-n}(d^{t}h_{n}).$
The left hand side of the equation becomes
$cg\cdot(-4\pi y)^{-(r+1)}\cdot\sum_{t=0}^{r}\binom{r}{t}(-1)^{t+1}\frac{\Gamma(\lambda+r)\Gamma(t+1)}{\Gamma(\lambda+t)}+g\cdot\delta_{\lambda}^{(r)}E.$
By comparing the coefficients of $(-4\pi y)^{t}$ for $t=0,1,\dots,r+1,$ one gets the desired equality. ∎
For a nomalized eigenform $f\in S_{k}(N,\chi)$, and a modular form $g\in M_{l}(N,\psi)$, define the completed Rankin–Selberg $L$-function (with a ‘classical’ normalization) as:
(4) $L(s,f\times g)=L_{\infty}(s,f\times g)L_{N}(2s+2-k-l,\chi\psi)D(s,f\times g),$
where $L_{\infty}(s,f\times g)=(2\pi)^{-2s}\Gamma(s)\Gamma(s+1-l)$ is the archimdean factor and for a Dirichlet character $\omega$ modulo $N$ put $L_{N}(s,\omega)=\sum_{n=1}^{\infty}\omega(n)n^{-s}$ with $\omega(n)=0$ if $(n,N)=1.$ If $g$ is also a normalized eigenform in $S_{l}(N,\psi),$ then $L(s,f\times g)$ satisfies a functional equation
(5) $L(s,f\times g)\approx L(k+l-1-s,f^{\rho},g^{\rho}).$
See Shimura [15, Section 3] and Hida [5, Section 9]. The line of symmetry for this functional equation is $\text{Re}(s)=(k+l-1)/2.$ An integer $m$ is critical for $L(s,f\times g),$ if the Gamma factors on both sides of the functional equation are finite at $s=m,$ i.e., if $\Gamma(m)\Gamma(m+1-l)$ and $\Gamma(k+l-1-m)\Gamma(k-m)$ are finite. Therefore the critical set is:
$\{m\in\mathbb{Z}\,|\,l\leq m\leq k-1\}.$
The following result of Shimura gives an algebraicity theorem for the values of the Rankin–Selberg convolution at critical points to the right of the line of symmetry:
Theorem 2.7.
Let $f$ be a normalized eigenform of $S_{k}(N)$, $g$ an element of $G_{l}(N)$, and $m$ a positive integer. Suppose $l<k$ and $\frac{1}{2}(k+l-2)<m<k$. Then $\pi^{-k}\langle f,f\rangle^{-1}D(m,f\times g)$ belongs to $\mathbb{Q}(f)\mathbb{Q}(g)$, the compositum of $\mathbb{Q}(f)$ and $\mathbb{Q}(g)$. Moreover, for every automorphism $\sigma$ of $\mathbb{C}$, we have
$\sigma\left(\pi^{-k}\langle f,f\rangle^{-1}D(m,f\times g)\right)\ =\ \pi^{-k}\langle f^{\sigma},f^{\sigma}\rangle^{-1}D(m,f^{\sigma}\times g^{\sigma}).$
One deduces that for an integer $m$ with $\frac{1}{2}(k+l-2)<m<m+1<k$ then
$\frac{D(m,f\times g)}{D(m+1,f\times g)}\in\mathbb{Q}(f)\mathbb{Q}(g).$
The parameters determining the abelian part and the infinite part of $L(s,f\times g)$ in (4) depend only on the weights and the nebentype characters involved. So for two $f,f^{\prime}\in S_{k}(N,\chi)$ and $g\in M_{l}(N,\psi)$ one has
(6) $L_{\infty}(s,f\times g)L_{N}(2s+2-k-l,\chi\psi)=L_{\infty}(s,f^{\prime}\times g)L_{N}(2s+2-k-l,\chi\psi).$
For brevity, let $\omega=\chi\psi.$ Suppose $\omega$ is a Dirichlet character modulo $N$ which is primitive. Then $L_{N}(s,\omega)=L(s,\omega)$. Let $\epsilon\in\{0,1\}$ such that $\omega(-1)=(-1)^{\epsilon}$. For $m\in\mathbb{N}$ and $m\equiv\epsilon\pmod{2}$ it is well known that
$L(m,\omega)=(-1)^{1+(m-\epsilon)/2}\frac{\mathfrak{g}(\omega)}{2\sqrt{-1}^{\epsilon}}\Bigl(\frac{2\pi}{N}\Bigr)^{m}\frac{B_{m,\bar{\omega}}}{m!},$
where $\mathfrak{g}(\omega)$ is the Gauss sum associated to $\omega$ and $B_{m,\bar{\omega}}$ is the generalized Bernoulli number for the character $\bar{\omega}=\omega^{-1}$. See Neukirch [12]. From this it follows that
(7) $\frac{L(m,\omega)}{L(m+2,\omega)}=(m+2)(m+1)\Bigl(\frac{N}{2\pi}\Bigr)^{2}\frac{B_{m,\bar{\omega}}}{B_{m+2,\bar{\omega}}}.$
If $\omega$ modulo $N$ is not primitive take $\omega^{\text{prim}}$ to be the primitive Dirichlet character defined by $\omega$. Then $L_{N}(s,\omega)=\prod_{p|N}\left(1-\omega^{\text{prim}}(p)p^{-s}\right)L(s,\omega^{\text{prim}}).$ Hence, from (7) and Theorem 2.7, for any integer $m$ with $\frac{k+l-2}{2}<m<m+1<k$, it follows that:
(8) $\frac{L(m,f\times g)}{L(m+1,f\times g)}\ \in\ \mathbb{Q}(f)\mathbb{Q}(g).$
This is proved in greater generality in the case of $L$-functions attached to cohomological cuspidal automorphic representations of $({\rm GL}_{n}\times{\rm GL}_{n^{\prime}})/F,$ when $nn^{\prime}$ is even and $F$ a totally real number field, by Harder and the second author [3].
2.2. An algorithm for the algebraic part of $D(m,f\times g)$
If $f\in S_{k}(\Gamma_{1}(N))$ is a newform then $f_{\rho}=\overline{f(-\overline{z})}=f^{\rho}$ is again a newform of $S_{k}(\Gamma_{1}(N)),$ where, recall that for any automorphism $\sigma$ of $\mathbb{C}$, one defines $f^{\sigma}:=\sum_{n=1}^{\infty}a(n,h)^{\sigma}q^{n}.$
Input: $f\in S_{k}(N,\chi)$ a newform of level $N,$ and $g\in S_{l}(N)$ an arbitrary cusp form with algebraic Fourier coefficients.
Output: The algebraic number $\frac{D(m,f\times g)}{\pi^{k}\langle f,f\rangle}$ for $\frac{k+l-2}{2}<m<k.$
  1. Step 1.
    Extend $f_{\rho}$ to a basis consisting of eigenforms $S_{k}(\Gamma_{1}(N))$ which we denote by $\mathcal{B}=\{f_{0}:=f_{\rho},f_{1},\dots,f_{n}\}$. (The dimension is $n$.)
  2. Step 2.
    Fix an $m$ in the critical set. Put r=k1mr=k-1-m. Pick the correct Eisenstein series $E_{\lambda,N}^{*}(z,\chi\psi)$ where $\lambda=k-l-2r$.
  3. Step 3.
    Find the holomorphic projection $h_{0}$ of $g\delta_{\lambda}^{(r)}E_{\lambda,N}(z,\chi\psi)$ using the recursive relations in Lemma 2.6.
  4. Step 4.
    Write the modular form $h_{0}$ in terms of $\mathcal{B}$, i.e., $h_{0}=\alpha_{0}f_{0}+\alpha_{1}f_{1}+\cdots+\alpha_{n}f_{n}$.
  5. Step 5.
    $\textbf{Return:}\quad c_{r}\alpha_{0}.$
This is the desired value due to Lemma 2.4, Lemma 2.3, Theorem 2.5, and $\langle f_{0},f_{i}\rangle=0$ for $i=1,\cdots,n$ and $\langle f^{\rho},f^{\rho}\rangle=\langle f,f\rangle$. The constant $c_{r}$ appears in Theorem 2.5.
This algorithm can be extended to the case when $f$ is only assumed to be an eigenform instead of newform. See Section 3.4 for an example.
2.3. An algorithm for the algebraic part of $L(m,f)$
We describe a well-known algorithm to calculate the algebraic part of the special value of $L(s,f)$ for the standard $L$-function of a new form $f$ at $s=m$ a critical value. The term algebraic is explained below. We need some preliminaries on computing $L$-values by modular symbols.
2.3.1. Special values of $L$-functions via modular symbols
Let $\Gamma$ be a congruence subgroup of $\mathrm{SL}_{2}(\mathbb{Z})$ and $\mathbb{P}^{1}(\mathbb{Q})=\mathbb{Q}\cup\{\infty\}.$ Define $\mathbb{M}_{2}$ to be the free abelian group on the symbols $\{\alpha,\beta\}\in\mathbb{P}^{1}(\mathbb{Q})$ modulo the relations $\{\alpha,\beta\}+\{\beta,\gamma\}+\{\gamma,\alpha\}=0$ for all $\alpha,\beta,\gamma\in\mathbb{P}^{1}(\mathbb{Q})$ and modulo all torsion. Let $\mathbb{Z}[X,Y]_{n}$ be the abelian group of homogeneous polynomials of degree $n$ in two variables $X$ and $Y$. Make $\mathbb{Z}[X,Y]_{n}$ into a $\Gamma$-module: if $\gamma=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\in\Gamma$ and $P\in\mathbb{Z}[X,Y]_{n}$ then define $(\gamma P)(X,Y)=P(dX-bY,-cX+aY).$ The abelian group $\mathbb{M}_{2}$ can also be made into a $\Gamma$-module by the action $g\{\alpha,\beta\}:=\{g\alpha,g\beta\}$ for all $g\in\Gamma$ and $\{\alpha,\beta\}\in\mathbb{M}_{2}$. Define the $\Gamma$-module $\mathbb{M}_{k}=\mathbb{Z}[X,Y]_{k-2}\otimes_{\mathbb{Z}}\mathbb{M}_{2}$ with the $\Gamma$ acting diagonally.
Definition 2.1.
(Modular Symbols) For an integer $k\geq 2$ and a congruence subgroup $\Gamma$, the space $\mathbb{M}_{k}(\Gamma)$ of weight $k$ modular symbols for $\Gamma$ is the quotient of $\mathbb{M}_{k}$ by all the relations $\gamma x-x$ for $x\in\mathbb{M}_{k}$, and $\gamma\in G$ and by any torsion.
For $P\in\mathbb{Z}[X,Y]_{k-2}$ and $\gamma\in\mathrm{SL}_{2}(\mathbb{Z})$ the associated Manin symbol is
$P,\gamma]:=\gamma(P\otimes\{0,\infty\})\in\mathbb{M}_{k}(\Gamma)$
When $\Gamma=\Gamma_{1}(N)$ and if $\gamma=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right),\ \gamma^{\prime}=\left(\begin{smallmatrix}a^{\prime}&b^{\prime}\\ c^{\prime}&d^{\prime}\end{smallmatrix}\right)\in\Gamma_{1}(N)$ are such that $(c,d)\equiv(c^{\prime},d^{\prime})\pmod{N}$, then $P,\gamma]=[P,\gamma^{\prime}]$. Hence, the Manin symbol $P,\gamma]$ is determined by $P$ and the lower two entries $c,d$ of the matrix $\gamma$. So we take $[P,(c,d)]$ itself to be a Manin symbol. For the following theorem, see, for example, Stein [17, Prop. 8.3].
Theorem 2.8.
The Manin Symbols generate $\mathbb{M}_{k}(\Gamma).$
One can define Hecke operators on the space of modular symbols; see Stein [17, Section 8.3].
Theorem 2.9.
Let $S_{k}(\Gamma),\bar{S}_{k}(\Gamma)$ denote the space of holomorphic and anti-holomorphic cusp forms respectively. Then the pairing $(\cdot,\cdot):(S_{k}(\Gamma)\oplus\bar{S}_{k}(\Gamma))\times\mathbb{M}_{k}(\Gamma)\rightarrow\mathbb{C}$
$((f_{1},f_{2}),P\otimes\{\alpha,\beta\})\ \mapsto\ \int_{\alpha}^{\beta}f_{1}(z)P(z,1)dz+\int_{\alpha}^{\beta}f_{2}(z)P(\bar{z},1)d\bar{z}$
is Hecke equivariant, i.e., for $x\in\mathbb{M}_{k}(\Gamma)$ we have $(T_{n}(f_{1},f_{2}),x)=(f_{1},f_{2},T_{n}(x)).$
For a cusp form $f\in S_{k}(N)$, the associated $L$-function is
(9) $D(s,f)=\sum_{n=1}^{\infty}\frac{a_{n}}{n^{s}},\quad{\rm and}\quad\int_{0}^{\infty}f(it)t^{s}\frac{dt}{t}=(2\pi)^{-s}\Gamma(s)D(s,f),\quad\Re(s)\gg 0.$
Consider the modular symbols $X^{j}Y^{k-2-j}\otimes\{0,\infty\}$ for $j=0,1,2,\dots,k-2$. Then the non-degenerate and Hecke equivariant bilinear pairing $(\cdot,\cdot)$ of these modular symbols against the cusp form $f$ is equal to
$(f,X^{j}Y^{k-2-j}\otimes\{0,\infty\})=\int_{0}^{i\infty}f(z)z^{j}dz=i^{j+1}\int_{0}^{\infty}f(it)t^{j}dt.$
Hence, by (9), we get the following relation between the bilinear pairing and the special value of $L$-function. For $m=1,2\dots,k-1$,
(10) $D(m,f)=\frac{(-2\pi\sqrt{-1})^{m}}{(m-1)!}\cdot(f,X^{m-1}Y^{k-2-(m-1)}\otimes\{0,\infty\}).$
For $m=1,2\dots,k-1$, define algebraic part of $L(m,f)$ by:
(11) $\frac{D(m,f)}{(-2\pi\sqrt{-1})^{m-1}D(1,f)}$ $=\frac{1}{(m-1)!}\cdot\frac{(f,X^{m-1}Y^{k-2-(m-1)}\otimes\{0,\infty\})}{(f,Y^{k-2}\otimes\{0,\infty\})}\;\;\text{for }m=1,3,\dots,$
(12) $\frac{D(m,f)}{(-2\pi\sqrt{-1})^{m-2}D(2,f)}$ $=\frac{1}{(m-1)!}\cdot\frac{(f,X^{m-1}Y^{k-2-(m-1)}\otimes\{0,\infty\})}{(f,X^{1}Y^{k-3}\otimes\{0,\infty\})}\;\;\text{for }m=2,4,\dots.$
2.3.2. The algorithm
Input: $f\in S_{k}(\Gamma_{1}(N))$ a newform. A prime $p$ for which $a(p,f)$ is also given.
Output: $\frac{D(m,f)}{(-2\pi\sqrt{-1})^{m-1}D(1,f)}$ for $m=1,3,\dots,$ and $\frac{D(m,f)}{(-2\pi\sqrt{-1})^{m-2}D(2,f)}$ for $m=2,4,\dots$.
  1. Step 1.
    For a given eigenform $f\in S_{k}(\Gamma_{1}(N))$ construct the space of modular symbols $\mathbb{M}_{k}(\Gamma_{1}(N)).$
  2. Step 2.
    Find a Manin basis for $\mathbb{M}_{k}(\Gamma_{1}(N))$. Call them $e_{1},e_{2},\dots,e_{n}$. Express the modular symbols $b_{i}:=X^{i-1}Y^{k-2-(i-1)}\otimes\{0,\infty\}$ for $i=1,2,\dots,k-1$ in terms of the basis $e_{1},e_{2},\dots,e_{n}$. Say $b_{i}=\sum_{j=1}^{n}a_{ji}e_{j}$ for $i=1,2,\dots,k-1$. Let $A$ be the matrix $a_{ij}]$.
  3. Step 3.
    For a prime $p$ find the matrix of the $p$-th Hecke operator $T_{p}$ with respect to the basis $e_{1},e_{2},\cdots e_{n}$. Call that matrix to be $B=[b_{ij}].$
  4. Step 4.
    Form a matrix $\mathcal{L}$ as follows:
    $\mathcal{L}=\left(\begin{array}[]{c|c}I_{k-1}&A^{t}\\ \hline\cr\\ \mathbf{0}_{n,k-1}&B-a(p,f)\cdot I_{n}\end{array}\right).$
  5. Step 5.
    Perform row reduction on $\mathcal{L}$. Call the resulting matrix $\mathcal{L}_{\text{red}}$. For the values $i=1,3,5\cdots$, find a column in $\mathcal{L}_{\text{red}}$ such that the first entry of the column is non-zero. Let such a column be $C_{\text{odd}}=(c_{1},c_{2},\cdots,c_{n+k-1})^{t}$. So $c_{1}\neq 0.$ Similarly, find a column $C_{\text{even}}=(d_{1},d_{2},\cdots,d_{n+k-1})^{t}$ of $\mathcal{L}_{\text{red}}$ such that $d_{2}\neq 0$.
  6. Step 6.
    Return: $\dfrac{1}{(i-1)!}\cdot\dfrac{c_{i}}{c_{1}}\text{ for $i=1,3,5,\cdots$ and }\dfrac{1}{(i-1)!}\cdot\dfrac{d_{i}}{d_{2}}\text{ for $i=2,4,\cdots$}.$
This is the desired value due to
$\frac{(f,X^{i-1}Y^{k-2-(i-1)}\otimes\{0,\infty\})}{(f,Y^{k-2}\otimes\{0,\infty\})}=\frac{c_{i}}{c_{1}}\>\>\>\mbox{for}\>\>i=1,3,5\cdots,$
$\frac{(f,X^{i-1}Y^{k-2-(i-1)}\otimes\{0,\infty\})}{(f,XY^{k-3}\otimes\{0,\infty\})}=\frac{d_{i}}{d_{2}}\>\>\>\mbox{for}\>\>i=2,4\cdots$
and the equations (11) and (12).
2.3.3. An example
Suppose $f$ is a newform of weight $5$ for $\Gamma_{1}(4)$ with $q$-expansion is $f(q)=q-4q^{2}+16q^{4}+\cdots$. The algorithm in 2.3.2 when used to compute the special values of $D(s,f)$ gives us:
$\frac{D(4,f)}{D(2,f)}=\frac{\pi^{2}}{5}\ \ \text{ and }\ \ \frac{D(3,f)}{D(1,f)}=\frac{5\pi^{2}}{12}.$
The interested reader is referred to the first author’s thesis [10] for details of the implementation of the algorithm.
2.4. Sturm’s criterion for congruence between cusp forms
To check congruence between cusp forms we will use the following criterion of Sturm [18]; see also Stein [17, p. 173].
Theorem 2.10 (Sturm).
Let $\mathfrak{P}$ be a prime ideal of a number field $K$. Let $f,g\in M_{k}(\Gamma,\mathcal{O}_{K})$ be modular forms of weight $k$ for a congruence subgroup $\Gamma$ of level $N$ with Fourier coefficients in the ring of integers $\mathcal{O}_{K}$ of $K$. Let $m=[SL_{2}(\mathbb{Z}):\Gamma].$ Suppose
$a(n,f)\equiv a(n,g)\pmod{\mathfrak{P}}$
for all
$n\leq\begin{cases}\frac{km}{12}-\frac{m-1}{N}&\mbox{if $f-g\in S_{k}(\Gamma,\mathcal{O}$}),\\ \frac{km}{12}&\mbox{otherwise.}\end{cases}$
Then $f\equiv g\pmod{\mathfrak{P}}$.
The bound for $n$ in the above theorem is called the Sturm bound for $\Gamma$ and weight $k$.
3. Examples
3.1. $S_{24}(\mathrm{SL}_{2}(\mathbb{Z}))\times S_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$
The dimension of $S_{24}(\mathrm{SL}_{2}(\mathbb{Z}))$ is $2$. There are no old forms as the level is $1$. The two distinct newforms, say, $f_{1}$} and $f_{2}$} have Fourier expansions:
$f_{1}$ $=q+(\beta_{0})q^{2}+\cdots,$
$f_{2}$ $=q+(-\beta_{0}+1080)q^{2}+\cdots,$
where $\beta_{0}=12\sqrt{144169}$. The coefficients are in the number field $\mathbb{Q}(\beta_{0}).$ The prime number $144169$ is ramified in the number field $\mathbb{Q}(\beta_{0})$. Let the prime ideal lying above $144169$ be $\mathfrak{P}_{144169}$. Also, $f_{1\rho}={f_{1}}^{\rho}={f_{1}}$ and $f_{2\rho}={f_{2}}^{\rho}={f_{2}}$ as the coefficients are totally real. Their difference is
$f_{1}-f_{2}=(2\beta_{0}-1080)q^{2}+\cdots.$
The ideal factorization of $(2\beta_{0}-1080)$ in the number field $\mathbb{Q}(\beta_{0})$ is
$(\left(2,\frac{1}{24}\beta_{0}-23\right))^{3}\cdot(\left(2,\frac{1}{24}\beta_{0}-22\right))^{3}\cdot(\left(3,\frac{1}{24}\beta_{0}-23\right))\\ \cdot(\left(3,\frac{1}{24}\beta_{0}-22\right))\cdot(\left(144169,\frac{1}{24}\beta_{0}+72062\right)).$
Therefore, since the Sturm bound is $2$ for $S_{24}(\mathrm{SL}_{2}(\mathbb{Z}))$, by Theorem 2.10 one has:
$f_{1}\equiv f_{2}\pmod{\mathfrak{P}_{144169}}$,
The space $S_{l}(\Gamma_{0}(N))=S_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$ is of dimension $1$ and generated by the unique Ramanujan cusp form $\Delta$. Put $g=\Delta.$ The Fourier expansion of $g$ is $q-24q^{2}+\cdots$.
Theorem 3.1.
For $f_{1},f_{2}\in S_{24}(\mathrm{SL}_{2}(\mathbb{Z}))$ and $g\in S_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$ as above, and any integer $m$ with $18\leq m\leq 22$, one has:
$\frac{{L}(m,f_{1}\times g)}{{L}(m+1,f_{1}\times g)}\equiv\frac{{L}(m,f_{2}\times g)}{{L}(m+1,f_{2}\times g)}\pmod{\mathfrak{P}_{144169}}$.
This is verified using the algorithm given in Section 2.2.
3.1.1. Values of $D(23,f_{1}\times g)$ and $D(23,f_{2}\times g)$
As Step 1, fix the basis $\mathcal{B}=\{f_{1},f_{2}\}$ of $S_{24}(\mathrm{SL}_{2}(\mathbb{Z}))$ consisting of newforms. Proceed to Step 2. The critical point is $23.$ So put $r=0$ and $\lambda=k-l-2r=12$. The Eisenstein series to be considered is $E_{\lambda,N}^{*}(z,1\!\!1)=E_{12,1}^{*}(z,1\!\!1).$ The space of Eisenstein series of weight $12$ and trivial character is of dimension $1$. The $q$-expansion of $E_{12,1}^{*}$ is
$E_{12,1}^{*}(z,1\!\!1)$ $=1+\frac{65520}{691}q+\frac{134250480}{691}q^{2}+\frac{11606736960}{691}q^{3}+O(q^{4}).$
Furthermore,
$g.\delta_{12}^{(0)}E_{12,1}^{*}(z,1\!\!1)=g.E_{12,1}^{*}(z,1\!\!1)=q+\frac{48936}{691}q^{2}+\frac{132852132}{691}q^{3}+O(q^{4}).$
Also, $g.E_{12,1}^{*}(z,1\!\!1)$ is a modular form (in fact, a cusp form) of weight $24$. So there is no need to take the holomorphic projection. Proceed to $\textit{Step 4}.$ In terms of the basis $\mathcal{B}=\{f_{1},f_{2}\}$
$g.E_{12,1}^{*}(z,1\!\!1)=\left(-\frac{27017}{2390898696}\beta_{0}+\frac{50418272}{99620779}\right)f_{1}+\left(\frac{27017}{2390898696}\beta_{0}+\frac{49202507}{99620779}\right)f_{2}.$
For Step 5 one gets
$\frac{D(23,f_{1}\times g)}{\pi^{24}\langle f_{1},f_{1}\rangle}=c_{0}\frac{\langle{f_{1}}^{\rho},g\delta_{\lambda}^{(r)}E_{\lambda,N}^{*}\rangle}{\pi^{24}\langle f_{1},f_{1}\rangle}=c_{0}\left(-\frac{27017}{2390898696}\beta_{0}+\frac{50418272}{99620779}\right).$
Similarly, for the pair $(f_{2},g)$ one gets
$\frac{D(23,f_{2}\times g)}{\pi^{24}\langle f_{2},f_{2}\rangle}=c_{0}\frac{\langle{f_{2}}^{\rho},g\delta_{\lambda}^{(r)}E_{\lambda,N}^{*}\rangle}{\pi^{24}\langle f_{2},f_{2}\rangle}=c_{0}\left(\frac{27017}{2390898696}\beta_{0}+\frac{49202507}{99620779}\right).$
3.1.2. Values of $D(22,f_{1}\times g)$ and $D(22,f_{2}\times g)$
As the basis is already fixed proceed to $\textit{Step 2}.$ For the critical point $22$ put $r=1$ and so $\lambda=k-l-2r=10$. So the Eisenstein Series is $E_{\lambda,N}^{*}(z,1\!\!1)=E_{10,1}^{*}(z,1\!\!1)$ and its $q$-expansion is
$E_{10,1}^{*}(z,1\!\!1)=1-264q-135432q^{2}-5196576q^{3}-69341448q^{4}-515625264q^{5}+O(q^{6}).$
Proceeding to Step 4, by Lemma 2.2, there are unique $h_{0}$ and $h_{1}$ sastisfying
$g\delta_{10}^{(1)}E_{10,1}^{*}(z,1\!\!1)=h_{0}+\delta_{22}^{(1)}h_{1}.$
Using the recursive relations in Lemma 2.6 calculate the holomorphic projection $h_{0}:$
$h_{0}$ $=-\frac{5}{11}q-\frac{24}{11}q^{2}-\frac{977148}{11}q^{3}+O(q^{4})$
$=\left(\frac{223}{38060616}\beta_{0}-\frac{365440}{1585859}\right)f_{1}+\left(-\frac{223}{38060616}\beta_{0}-\frac{355405}{1585859}\right)f_{2}.$
Proceeding to the last Step 6,
$\frac{D(22,f_{1}\times g)}{\pi^{24}\langle f_{1},f_{1}\rangle}=\frac{\langle f_{1}^{\rho},\delta_{10}^{(1)}E_{10,1}^{*}\rangle}{\pi^{24}\langle f_{1},f_{1}\rangle}=\frac{\langle f_{1},h_{0}\rangle}{\pi^{24}\langle f_{1},f_{1}\rangle}=c_{1}\left(\frac{223}{38060616}\beta_{0}-\frac{365440}{1585859}\right).$
Similarly,
$\frac{D(22,f_{2}\times g)}{\pi^{24}\langle f_{2},f_{2}\rangle}=c_{1}\left(-\frac{223}{38060616}\beta_{0}-\frac{355405}{1585859}\right).$
3.1.3. Ratios of successive critical values
Since $N=1$ and $\chi$ and $\psi$ are trivial, we have $L_{N}(s,\chi\psi)=\zeta(s)$, the Riemann Zeta function. It is well known that:
$\zeta(2\cdot 23-k-l)=\zeta(10)=\frac{691}{638512875}\,\pi^{12},\quad\zeta(2\cdot 22+2-k-l)=\zeta(8)=\frac{1}{93555}\,\pi^{10}.$
Therefore, for $i=1,2$
$\frac{L_{\infty}(22,f_{i}\times g)\zeta(8)}{L_{\infty}(23,f_{i}\times g)\zeta(10)}=\frac{13650}{83611}.$
The ratios of the special values of the completed $L$-functions are
$\frac{L(22,f_{1}\times g)}{L(23,f_{1}\times g)}=-\frac{1}{1905750}\beta_{0}+\frac{51866}{317625}\quad\text{ and }\quad\frac{L(22,f_{1}\times g)}{L(23,f_{2}\times g)}=\frac{1}{1905750}\beta_{0}+\frac{51686}{317625}.$
The ideal factorization of the quantity $\frac{L(22,f_{1}\times g)}{L(23,f_{1}\times g)}-\frac{L(22,f_{2}\times g)}{L(23,f_{2}\times g)}$ in the number field $\mathbb{Q}(\beta_{0})$ is
$(\left(2,\frac{1}{24}\beta_{0}-23\right))^{2}\cdot(\left(2,\frac{1}{24}\beta_{0}-22\right))^{2}\cdot(\left(3,\frac{1}{24}\beta_{0}-23\right))^{-1}\cdot(\left(3,\frac{1}{24}\beta_{0}-22\right))^{-1}\cdot\\ (\left(5,\frac{1}{24}\beta_{0}-21\right))^{-3}\cdot(\left(5,\frac{1}{24}\beta_{0}-19\right))^{-3}\cdot(\left(7,\frac{1}{24}\beta_{0}-20\right))^{-1}\cdot(\left(7,\frac{1}{24}\beta_{0}-18\right))^{-1}\cdot\\ (\left(11,\frac{1}{24}\beta_{0}-20\right))^{-2}\cdot(\left(11,\frac{1}{24}\beta_{0}-14\right))^{-2}\cdot(\left(144169,\frac{1}{24}\beta_{0}+72062\right)).$
Hence
$\frac{L(22,f_{1}\times g)}{L(23,f_{1}\times g)}\equiv\frac{L(22,f_{2}\times g)}{L(23,f_{2}\times g)}\pmod{\mathfrak{P}_{144169}}$.
3.1.4. Other ratios
For $m=21,20,19$, the ideal factorizations of
$\frac{{L}(m,f_{1}\times g)}{{L}(m+1,f_{1}\times g)}-\frac{{L}(m,f_{2}\times g)}{{L}(m+1,f_{2}\times g)}$
in the number field $\mathbb{Q}(\beta_{0})$ are stated below. The method of computation is along the same lines as in the previous sub-sections.
$\frac{L(21,f_{1}\times g)}{L(22,f_{1}\times g)}-\frac{L(21,f_{2}\times g)}{L(22,f_{2}\times g)}\\ =(\left(2,\frac{1}{24}\beta_{0}-23\right))^{-1}\cdot(\left(2,\frac{1}{24}\beta_{0}-22\right))^{-1}\cdot(\left(3,\frac{1}{24}\beta_{0}-23\right))^{-3}\cdot(\left(3,\frac{1}{24}\beta_{0}-22\right))^{-3}\cdot\\ (\left(5,\frac{1}{24}\beta_{0}-21\right))^{-1}\cdot(\left(5,\frac{1}{24}\beta_{0}-19\right))^{-1}\cdot(\left(7,\frac{1}{24}\beta_{0}-20\right))^{-1}\cdot(\left(7,\frac{1}{24}\beta_{0}-18\right))^{-1}\cdot\\ (\left(17,\frac{1}{24}\beta_{0}-21\right))^{-1}\cdot(\left(17,\frac{1}{24}\beta_{0}-7\right))^{-1}\cdot(\left(144169,\frac{1}{24}\beta_{0}+72062\right)),$
$\frac{L(20,f_{1}\times g)}{L(21,f_{1}\times g)}-\frac{L(20,f_{2}\times g)}{L(21,f_{2}\times g)}\\ =(\left(2,\frac{1}{24}\beta_{0}-23\right))^{-3}\cdot(\left(2,\frac{1}{24}\beta_{0}-22\right))^{-3}\cdot(\left(3,\frac{1}{24}\beta_{0}-23\right))^{-1}\cdot(\left(3,\frac{1}{24}\beta_{0}-22\right))^{-1}\cdot\\ (\left(5,\frac{1}{24}\beta_{0}-21\right))^{-1}\cdot(\left(5,\frac{1}{24}\beta_{0}-19\right))^{-1}\cdot(\left(103,\frac{1}{24}\beta_{0}+18\right))^{-1}\cdot(\left(103,\frac{1}{24}\beta_{0}+40\right))^{-1}\cdot\\ (\left(144169,\frac{1}{24}\beta_{0}+72062\right)),$
and,
$\frac{L(19,f_{1}\times g)}{L(20,f_{1}\times g)}-\frac{L(19,f_{2}\times g)}{L(20,f_{2}\times g)}\\ =(\left(2,\frac{1}{24}\beta_{0}-23\right))^{-1}\cdot(\left(2,\frac{1}{24}\beta_{0}-22\right))^{-1}\cdot(\left(7,\frac{1}{24}\beta_{0}-20\right))^{-1}\cdot(\left(7,\frac{1}{24}\beta_{0}-18\right))^{-1}\cdot\\ (\left(17,\frac{1}{24}\beta_{0}-21\right))^{-1}\cdot(\left(17,\frac{1}{24}\beta_{0}-7\right))^{-1}\cdot(\left(19,\frac{1}{24}\beta_{0}-15\right))^{-1}\cdot(\left(19,\frac{1}{24}\beta_{0}-11\right))^{-1}\cdot\\ (\left(144169,\frac{1}{24}\beta_{0}+72062\right)).$
3.1.5. The critical point $m=18$
When the critical point under consideration is $m=18,$ the Eisenstein series is no longer holomorphic. So $r=5$ and $\lambda=24-12-2r=2$; the characters of $f,f^{\prime}$ and $g$ are all trivial. The correct Eisenstein series is a non-holomorphic one:
$E_{2,1}^{*}(z,1\!\!1)=\frac{-12}{4\pi y}+1-24q-72q^{2}$-$\cdots$
See Miyake [9, Chapter 7]. By using Lemma 2.6 the holomorphic projection $h_{0}$ to be
$h_{0}$ $=-\frac{5}{24871}q+\frac{10536}{24871}q^{2}-\frac{212004}{3553}q^{3}+O(q^{4})$
$=\left(\frac{1103}{86055052776}\beta_{0}-\frac{385240}{3585627199}\right)f_{1}+\left(-\frac{1103}{86055052776}\beta_{0}-\frac{335605}{3585627199}\right)f_{2}.$
A similar calculation like in the previous sub-sections yield the ideal factorization of
$\frac{L(18,f_{1}\times g)}{{L}(19,f_{1}\times g)}-\frac{L(18,f_{2}\times g)}{L(19,f_{2}\times g)}\\ =(\left(3,\frac{1}{24}\beta_{0}-23\right))^{-1}\cdot(\left(3,\frac{1}{24}\beta_{0}-22\right))^{-1}\cdot(\left(5,\frac{1}{24}\beta_{0}-21\right))^{-1}\cdot(\left(5,\frac{1}{24}\beta_{0}-19\right))^{-1}\cdot\\ (\left(7,\frac{1}{24}\beta_{0}-20\right))^{-1}\cdot(\left(7,\frac{1}{24}\beta_{0}-18\right))^{-1}\cdot(\left(17,\frac{1}{24}\beta_{0}-21\right))^{-1}\cdot(\left(17,\frac{1}{24}\beta_{0}-7\right))^{-1}\cdot\\ (\left(144169,\frac{1}{24}\beta_{0}+72062\right)).$
This completes the verification of Thm. 3.1 for all successive ratios of critical points to the right of the line of symmetry.
3.2. $S_{30}(\mathrm{SL}_{2}(\mathbb{Z}))\times S_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$
The space $S_{30}(\mathrm{SL}_{2}(\mathbb{Z}))$ is again of dimension $2$. Call the newforms $f_{1}$ and $f_{2}$. The form $f_{1}$ is a Galois conjugte of $f_{2}$. The Fourier coefficients of $f_{1}$ and $f_{2}$ lie in the number field $\mathbb{Q}(\beta_{1})$ where $\beta_{1}$ satisfies the polynomial $x^{2}-8640x-454569984.$ The prime $51349$ is ramified in $\mathbb{Q}(\beta_{1})$. Let $\mathfrak{P}_{51349}$ be the prime ideal lying above $51349$. Like in 3.1, one verifies that $f_{1}\equiv f_{2}\pmod{\mathfrak{P}_{51349}}$. Take $g=\Delta$ to be Ramanuajan cusp form which is of weight $12$ and level $1$. Then exactly like in the 3.1 we get:
Theorem 3.2.
For $f_{1},f_{2}\in S_{30}(\mathrm{SL}_{2}(\mathbb{Z}))$ and $g\in S_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$ as above, and for an integer $m$ with $21\leq m\leq 28$, one has:
$\frac{L(m,f_{1}\times g)}{L(m+1,f_{1}\times g)}\equiv\frac{L(m,f_{2}\times g)}{L(m+1,f_{2}\times g)}\pmod{\mathfrak{P}_{51349}}$.
$3.$ $S_{13}(\Gamma_{0}(3),\chi)\times S_{6}(\Gamma_{0}(3))$
Let $\chi$ be the unique quadratic character modulo $3$. All the eigenforms in $S_{13}(\Gamma_{1}(3))=S_{13}(\Gamma_{0}(3),\chi)$ are newforms. The $q$-expansions of the newforms are
$f_{1}$ $=q+729q^{3}+(4096)q^{4}+O(q^{5}),$
$f_{2}$ $=q+\nu_{1}q^{2}+(-3\nu_{1}-675)q^{3}+(-4328)q^{4}+O(q^{5}),$
$f_{3}$ $=q+-\nu_{1}q^{2}+(3\nu_{1}-675)q^{3}+(-4328)q^{4}+O(q^{5}),$
where $\nu_{1}=\sqrt{-8424}$. It can checked that $f_{2}$ and $f_{3}$ are Galois conjugates of each other, and in fact, $f_{2\rho}=f_{2}^{\rho}=f_{3}$ and $f_{3\rho}=f_{3}^{\rho}=f_{2}$ as the coefficients are in the imaginary quadratic field $\mathbb{Q}(\nu_{1}).$ However, $f_{1}$ and $f_{2}$ are not Galois conjugate of each other, since the coefficients of $f_{1}$ are in $\mathbb{Q}$; also, $f_{1\rho}=f_{1}^{\rho}=f_{1},$ Using the Sturm bound (Thm. 2.10) one verifies
$f_{1}\equiv f_{2}\pmod{\mathfrak{P}_{13}}$,
where $\mathfrak{P}_{13}=(\left(13,\frac{1}{18}\nu_{1}\right))$ is the prime ideal lying above $(13)\subset\mathbb{Z}$ in the number field $\mathbb{Q}(\nu_{1}).$
The space $S_{6}(\Gamma_{0}(3))$ is one-dimensional, spanned by $g$ whose fourier expansion is
$g=q-6q^{2}+9q^{3}+4q^{4}+O(q^{5}).$
3.3.1. Ratios of successive critical values
Theorem 3.3.
For $f_{1},f_{2}\in S_{13}(\Gamma_{0}(3),\chi)$ and $g\in S_{6}(\Gamma_{0}(3))$ as above, and for an integer $m$ with $9\leq m\leq 11$, one has:
$\frac{L(m,f_{1}\times g)}{L(m+1,f_{1}\times g)}\equiv\frac{L(m,f_{2}\times g)}{L(m+1,f_{2}\times g)}\pmod{\mathfrak{P}_{13}}$.
The ratios of the Rankin-Selberg $L$-values are calculated exactly as in the previous subsections. The reader should keep in mind that $f_{2}^{\rho}=f_{3},f_{3}^{\rho}=f_{2}$, $\langle f_{3},f_{3}\rangle=\langle f_{3}^{\rho},f_{3}^{\rho}\rangle=\langle f_{2},f_{2}\rangle$, and $L_{N}(s,\chi\psi)=L(s,\psi)$ is the Dirichlet $L$-function for the unique quadratic character modulo $3$. The ideal factorization of the difference of the ratios $\frac{L(m,f_{1}\times g)}{L(m+1,f_{1}\times g)}-\frac{L(m,f_{2}\times g)}{L(m+1,f_{2}\times g)}$, for $m=11$, in the number field $\mathbb{Q}(\nu_{1})$ is given by:
$\frac{L(11,f_{1}\times g)}{L(12,f_{1}\times g)}-\frac{L(11,f_{2}\times g)}{L(12,f_{2}\times g)}=(\left(2,\frac{1}{18}\nu_{1}\right))^{2}\cdot(\left(3,\frac{1}{18}\nu_{1}+1\right))^{2}\cdot(\left(3,\frac{1}{18}\nu_{1}+2\right))\\ \cdot(\left(5,\frac{1}{18}\nu_{1}+2\right))^{-1}\cdot(\left(5,\frac{1}{18}\nu_{1}+3\right))^{-1}\cdot\left(11\right)^{-1}\cdot(\left(13,\frac{1}{18}\nu_{1}\right))\cdot(\left(71,\frac{1}{18}\nu_{1}+20\right))^{-1}.$
So they are congruent modulo $\mathfrak{P}_{13}$; the same conclusion holds for the other ratios as well.
3.4. $S_{26}(\mathrm{SL}_{2}(\mathbb{Z}))\times S_{13}(\Gamma_{0}(3),\psi)$
Unlike the previous examples, here the cusp form of higher weight $k$ is fixed, and the modular forms of lower weight $l$ vary in a congruence class.
For the weight $k=26$ cusp form, take $f\in S_{26}(\mathrm{SL}_{2}(\mathbb{Z}))$ with $q$-expansion
$f=q-48q^{2}-195804q^{3}-33552128q^{4}-741989850q^{5}+9398592q^{6}+O(q^{7}).$
and view it as an element of $S_{26}(\Gamma_{1}(3)).$
We take two cusp forms of weight $13$ and level $3$. In order to have consistency of notations call the forms $f_{1}$ and $f_{2}$ from 3.3 as $g_{1}$ and $g_{2}$ in $S_{13}(\Gamma_{0}(3),\psi)$, respectively, where $\psi$ is the non-trivial primitive character modulo $3$. It is already established that $g_{1}\equiv g_{2}\pmod{\mathfrak{P}_{13}}$, where $\mathfrak{P}_{13}$ is the prime ideal lying above $(13)\subset\mathbb{Z}$ in $\mathbb{Q}(\nu_{1}),$ where $\nu_{1}^{2}=-8424$.
As a Hecke module, the $7$-dimensional space $S_{26}(\Gamma_{1}(3))$ decomposes into a sum of:
  1. (1)
    a $2$-dimensional subspace with a basis of the old forms $f$ and $\hat{f}:=\langle 3\rangle\cdot f=f(3z),$
  2. (2)
    a $2$-dimensional subspace with a basis consisting of a newform form and its nontrivial Galois conjugate,
  3. (3)
    a $3$-dimensional subspace with a basis consisting of a newform form and its two distinct Galois conjugates.
This data is obtained from the $L$-functions and modular forms database LMFDB [7]. As there are in total $5$ newforms, finding an extension of $\mathbb{Q}(\nu_{1})$ containing all the Fourier coefficients of such newforms is computationally taxing. Instead, as a work around, we compute a basis consisting of Fourier coefficients in $\mathbb{Q},$ for the space in $(2)$ and for the space in $(3).$ This can be achieved in SAGE by applying the decomposition() command on any Hecke operator away from the level.
If $h_{0}$ is the holomorphic projection and $h_{0}=c_{0}f+c_{1}\hat{f}+\cdots\in S_{26}(\Gamma_{1}(3))$ then it is no longer true that $\langle f,h_{0}\rangle=c_{0}\langle f,f\rangle,$ since $\langle f,\hat{f}\rangle\neq 0.$ Hence $c_{r}\alpha_{0}$ is not the correct $L$-value. So while returning the $L$-value this has to be taken into account. The relation between $\langle f,f\rangle$ and $\langle f,\hat{f}\rangle$ is as follows: Suppose $f\in S_{k}(\Gamma_{1}(N))$ be a newform. Let $a(p,f)$ be its coeffecient at $p$. In $S_{k}(\Gamma_{0}(Np)),$ taking $\hat{f}(z)=f(pz)$ one has:
$\langle\hat{f},f\rangle_{\Gamma_{0}(Np)}=\langle f(pz),f(z)\rangle_{\Gamma_{0}(Np)}=\langle f(z),f(pz)\rangle_{\Gamma_{0}(Np)}=\frac{p^{-k+1}a(p,f)}{1+p}\langle f,f\rangle_{\Gamma_{0}(Np)}.$
See Bellaïche [6, p. 284].
3.4.1. Ratios of special values
In this example, we verify the congruence in all cases except one recorded in the following theorem:
Theorem 3.4.
For $f\in S_{26}(\mathrm{SL}_{2}(\mathbb{Z}))$ and $g_{1},g_{2}\in S_{13}(\Gamma_{0}(3),\psi)$ as above, and an integer $m$ with $19\leq m\leq 23$, one has:
$\frac{L(m,f\times g_{1})}{L(m+1,f\times g_{1})}\equiv\frac{L(m,f\times g_{2})}{L(m+1,f\times g_{2})}\pmod{\mathfrak{P}_{13}}$.
For the ratios of the right extreme critical values, one has for the Rankin–Selberg convolutions, the congruence:
$\frac{D(24,f\times g_{1})}{D(25,f\times g_{1})}\,\,\,{\equiv}\,\,\,\frac{D(24,f\times g_{2})}{D(25,f\times g_{2})}\pmod{\mathfrak{P}_{13}}$,
however, for the ratios of completed $L$-values, one has:
$\frac{L(24,f\times g_{1})}{L(25,f\times g_{1})}\,\,\,{\not\equiv}\,\,\,\frac{L(24,f\times g_{2})}{L(25,f\times g_{2})}\pmod{\mathfrak{P}_{13}}$.
For $19\leq m\leq 23$, the congruence is verified as in the previous sections. In the exceptional case $m=24$, the non-congruence for the ratios of extreme critical values is explained like this. The character of $f\in S_{k}(\Gamma_{1}(3))$ is trivial modulo $3$ and the character of $g_{1},g_{2}\in S_{k}(\Gamma_{0}(3),\psi)$ is the non-trivial quadratic character modulo $3$. So
$\frac{L_{\infty}(24,f\times g_{1})L_{N}(2\cdot 24+2-26-13,\chi\psi)}{L_{\infty}(25,f\times g_{1})L_{N}(2\cdot 25+2-26-13,\chi\psi)}\ =\ \frac{L_{\infty}(24,f\times g_{2})L(11,\psi)}{L_{\infty}(25,f\times g_{2})L(13,\psi)}=\frac{60951}{444808}.$
The first few prime ideals (written in the increasing order of their norms) appearing in the ideal factorization of the quantity $\frac{L(24,f\times g_{1})}{L(25,f\times g_{1})}-\frac{L(24,f\times g_{2})}{L(25,f\times g_{2})}$ in the number field $\mathbb{Q}(\nu_{1})$ is
$(\left(2,\frac{1}{18}\nu_{1}\right))^{-9}\cdot(\left(3,\frac{1}{18}\nu_{1}+1\right))^{2}\cdot(\left(5,\frac{1}{18}\nu_{1}+2\right))^{-1}\cdot(\left(5,\frac{1}{18}\nu_{1}+3\right))^{-1}\cdot(\left(7,\frac{1}{18}\nu_{1}+3\right))^{-2}\\ \cdot(\left(7,\frac{1}{18}\nu_{1}+4\right))^{-2}\cdot(\left(13,\frac{1}{18}\nu_{1}\right))^{-2}\cdot\left(19\right)^{-1}\cdot(\left(31,\frac{1}{18}\nu_{1}+6\right))^{-1}\cdots.$
Hence
(13) $\frac{L(24,f\times g_{1})}{L(25,f\times g_{1})}\ {\not\equiv}\ \frac{L(24,f\times g_{2})}{L(25,f\times g_{2})}\pmod{\mathfrak{P}_{13}}$.
In fact,
(14) $v_{\mathfrak{P}_{13}}\left(\frac{D(24,f\times g_{1})}{D(25,f\times g_{1})}-\frac{D(24,f\times g_{2})}{D(25,f\times g_{2})}\right)=1.$
The reason for non-congruence is due to the ratio $\frac{L(11,\psi)}{L(13,\psi)}$ of the abelian $L$-functions. Since $L(13,\psi)$ has in it the generalized Bernoulli number $B_{13,\psi}=-1445626/3=-1\cdot 2\cdot 3^{-1}\cdot 7\cdot 13^{3}\cdot 47,$ the congruence modulo $\mathfrak{P}_{13}$ in (14) gets ‘cancelled’ by the $13^{3}$ in the denominator of $\frac{L(11,\psi)}{L(13,\psi)}$ explaining the non-congruence in (13).
3.5. $S_{24}(\mathrm{SL}_{2}(\mathbb{Z}))\times M_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$
We may also vary the modular forms in a congruence class with one being cuspidal and the other an Eisenstein series. In this section, we vary the lower weight modular forms of weight $l=12$, with one form being the Ramanujan $\Delta$-function, and the other form the weight $12$ Eisenstein series. More precisely, let $f$ be the newform of weight $k=24$ for $SL_{2}(\mathbb{Z})$ with $q$-expansion $f(q)=q+\beta_{0}q^{2}+\left(-48\beta_{0}+195660\right)q^{3}+O(q^{4}),$ where $\beta_{0}$ satisfies the polynomial $x^{2}-1080x-20468736\in\mathbb{Q}[x]$. This is one of the newforms from Section 3.1. For the smaller weight $l=12$, take $g_{1}:=\Delta\in S_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$ and $g_{2}:=E_{12}(z)={691/65520}+q+\cdots\in M_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$, the Eisenstein series of weight $12$ and level $1$. The well-known Ramanujan congruence asserts that $g_{1}\equiv g_{2}\pmod{691}$.
Theorem 3.5.
For $f\in S_{24}(\mathrm{SL}_{2}(\mathbb{Z}))$ and $g_{1},g_{2}\in M_{12}(\mathrm{SL}_{2}(\mathbb{Z}))$ as above
$\frac{L(m,f\times g_{1})}{L(m+1,f\times g_{1})}\ \equiv\ \frac{L(m,f\times g_{2})}{L(m+1,f\times g_{2})}\pmod{\mathfrak{P}_{691}}\>\>\>\mbox{for}\>\>m=22,21,20,19,18.$
The $p$-th Fourier coefficient of $E_{12}$ being $p^{11}+1$, one has $L(s,f\times E_{12})=L(s,f)L(s+11,f)$, it is probably more enlightening to write the above congruence as:
$\frac{L(m,f\times\Delta)}{L(m+1,f\times\Delta)}\ \equiv\ \frac{L(m,f)L(m-11,f)}{L(m+1,f)L(m-10,f)}\pmod{\mathfrak{P}_{691}}$
In a similar vein, by Shimura [15, Lem. 1], one has:
$D(s,f\times g_{2})=\frac{D(s,f)\cdot D(s-11,f)}{\zeta(2s-34)},$
where $\zeta(s)$ is the Riemann zeta function and $D(s,f)=\sum_{n=1}^{\infty}a(n,f)n^{-s}$. Therefore,
$\dfrac{D(m,f,g_{2})}{D(m+1,f,g_{2})}$ $=\frac{\zeta(2m+2-34)}{\zeta(2m-34)}\dfrac{D(m,f)}{D(m-10,f)}\dfrac{D(m-11,f)}{D(m+1,f)}$
$=\left(\frac{1}{(2\pi\sqrt{-1})^{2}}\frac{\zeta(2m+2-34)}{\zeta(2m-34)}\right)\cdot\left(\dfrac{D(m,f)}{(2\pi\sqrt{-1})^{-1}D(m-10,f)}\right)$
$\qquad\qquad\qquad\cdot\left(\dfrac{D(m-11,f)}{(2\pi\sqrt{-1})^{-1}D(m+1,f)}\right).$
The first term inside $\left(\cdots\right)$ is rational due to well known result on the special values of the Riemann zeta function. The second and the third term are algebraic due to equations (11) and (12). Hence we use the algorithm described in Section 2.3.2 to calculate the ratios of the special values $\frac{L(m,f\times g_{2})}{L(m+1,f\times g_{2})}.$ The ratios $\frac{D(m,f)}{(-2\pi\sqrt{-1})^{m-1}D(1,f)}$ and $\frac{D(m,f)}{(-2\pi\sqrt{-1})^{m-2}D(2,f)}$ depending on the pairity of $m$ are summarized in the table.
$m$ $\frac{D(m,f)}{(-2\pi\sqrt{-1})^{m-1}D(1,f)}$
11 11
33 $\frac{569}{18825760800}\beta_{0}-\frac{16757416}{196101675}$
55 $\frac{3403}{218993544000}\beta_{0}+\frac{40511069}{4562365500}$
77 $\frac{147089}{17869873190400}\beta_{0}-\frac{854671739}{744578049600}$
99 $\frac{6538127}{1250891123328000}\beta_{0}+\frac{10060850717}{52120463472000}$
$11$ $\frac{320477}{77436117158400}\beta_{0}-\frac{169463087}{3226504881600}$
$13$ $\frac{320477}{77436117158400}\beta_{0}+\frac{169463087}{3226504881600}$
$15$ $\frac{6538127}{1250891123328000}\beta_{0}-\frac{10060850717}{52120463472000}$
$17$ $\frac{147089}{17869873190400}\beta_{0}+\frac{854671739}{744578049600}$
1919 $\frac{3403}{218993544000}\beta_{0}-\frac{40511069}{4562365500}$
2121 $\frac{569}{18825760800}\beta_{0}+\frac{16757416}{196101675}$
$23$ 1-1
$m$ $\frac{D(m,f)}{(-2\pi\sqrt{-1})^{m-2}D(2,f)}$
22 11
44 $\frac{1}{15120000}\beta_{0}-\frac{59221}{630000}$
66 $\frac{1}{26593920}\beta_{0}+\frac{12031}{1108080}$
88 $\frac{1783}{80372736000}\beta_{0}-\frac{5312293}{3348864000}$
$10$ $\frac{37}{2344204800}\beta_{0}+\frac{31327}{97675200}$
$12$ $\frac{847}{60279552000}\beta_{0}-\frac{371437}{2511648000}$
$14$ $\frac{37}{2344204800}\beta_{0}+\frac{31327}{97675200}$
$16$ $\frac{1783}{80372736000}\beta_{0}-\frac{5312293}{3348864000}$
$18$ $\frac{1}{26593920}\beta_{0}+\frac{12031}{1108080}$
2020 $\frac{1}{15120000}\beta_{0}-\frac{59221}{630000}$
$22$ 11
From the table above one concludes
$\frac{L(22,f\times g_{2})}{L(23,f\times g_{2})}$ $=-\frac{19153}{3052249200}\beta_{0}+\frac{23359724}{63588525}.$
Similarly, for the other ratios in the right hand side of Thm. 3.5 are determined. For the left hand side of Thm. 3.5, note that $\frac{D(m,f\times g_{1})}{\pi^{24}\langle f\times f\rangle}$ and the ratios of successive critical values $\frac{L(m,f\times g_{1})}{L(m+1,f\times g_{1})}$ have already been calculated in Sect. 3.1. The prime ideal factorization of the difference of the ratios has the following shape:
$\frac{L(22,f\times g_{1})}{L(23,f\times g_{1})}-\frac{L(22,f\times g_{2})}{L(23,f\times g_{2})}=\cdots(\left(13,\frac{1}{24}\beta_{0}-20\right))^{-1}\cdot(\left(73,\frac{1}{24}\beta_{0}-6\right))\\ \cdot(\left(691,\frac{1}{24}\beta_{0}-12\right))\cdot(\left(691,\frac{1}{24}\beta_{0}+658\right))\cdot(\left(23003,\frac{1}{24}\beta_{0}+2325\right)).$
The only prime that ramifies in $\mathbb{Q}(\beta_{0})$ is $144169.$ The ideal $\mathfrak{P}_{691}:=(691)\subset\mathbb{Q}(\beta_{0})$ is split and has the factorization $(\left(691,\frac{1}{24}\beta_{0}-12\right))\cdot(\left(691,\frac{1}{24}\beta_{0}+658\right))$. Hence
$\frac{L(22,f\times g_{1})}{L(23,f\times g_{1})}\equiv\frac{L(22,f\times g_{2})}{L(23,f\times g_{2})}\pmod{\mathfrak{P}_{691}}$.
Same conclusion holds for other ratios as well.
4. A conjecture on the ratios of the special values of the Rankin-Selberg L-functions
The main purpose of this article is to investigate the principle:
$f\equiv f^{\prime}\pmod{\mathfrak{P}}\ \ \implies\ \ \frac{L(m,f\times g)}{L(m+1,f\times g)}\equiv\frac{L(m,f^{\prime}\times g)}{L(m+1,f^{\prime}\times g)}\pmod{\mathfrak{P}}$.
Keeping the above examples in mind which lend credence to this principle, while accounting for the exceptional situation as in Thm. 3.4, we formulate the following conjecture:
Conjecture 4.1.
Let $f,f^{\prime}\in S_{k}(N,\chi)$ and $g\in S_{l}(M,\psi)$ be normalized eigenforms with kl>2k-l>2. Let $\mathfrak{P}\subset\mathbb{Q}(f)\mathbb{Q}(f^{\prime})\mathbb{Q}(g)$ be a prime ideal such that
$a(n,f)\equiv a(n,f^{\prime})\pmod{\mathfrak{P}}\;\;\text{ for all }n\in\mathbb{N}.$
For an integer $m$ with $\frac{k+l}{2}-1<m\leq k-2,$ assume:
(15) $v_{\mathfrak{P}}\left(\frac{L_{\infty}(m,f\times g)L_{N}(2m+2-k-l,\chi\psi)}{L_{\infty}(m+1,f\times g)L_{N}(m+4-k-l,\chi\psi)}\right)\geq 0.$
Then for all such $m$ one has the congruence:
(16) $\frac{L(m,f\times g)}{L(m+1,f\times g)}\ \equiv\ \frac{L(m,f^{\prime}\times g)}{L(m+1,f^{\prime}\times g)}\pmod{\mathfrak{P}}$.
Now, let $f\in S_{k}(N,\chi)$ and $g,g^{\prime}\in S_{l}(M,\psi)$ be normalized eigenforms and $\mathfrak{P}\subset\mathbb{Q}(f)\mathbb{Q}(g^{\prime})\mathbb{Q}(g)$ be a prime ideal such that
$a(n,g)\equiv a(n,g^{\prime})\pmod{\mathfrak{P}}\text{ for all }n\in\mathbb{N}.$
Assume (15) is satisfied. Then
(17) $\frac{L(m,f\times g)}{L(m+1,f\times g)}\ \equiv\ \frac{L(m,f\times g^{\prime})}{L(m+1,f\times g^{\prime})}\pmod{\mathfrak{P}}$.
In the preceding section, we computationally verified this Conjecture 4.1 in many examples. Especially, note that the hypothesis (15) has been introduced to take care of the exceptional situation in Thm. 3.4. As mentioned in the introduction, in the companion article [11], we prove a variation of this conjecture using the machinery of Eisenstein cohomology as developed by Harder and the second author in [3].
Acknowledgements: The authors are grateful to the Institute for Advanced Study, Princeton, for a summer collaborator’s grant in 2023 when this project got started. The authors also thank Baskar Balasubramanyam for some invaluable comments. The calculations were carried out at the Bhaskara Lab of IISER Pune. The first author is supported by the CSIR fellowship for his Ph.D.